OpenStudy (anonymous):

factor completely x2 + 9x + 20

Parth (parthkohli):

\( \color{Black}{\Rightarrow x^2 + 4x + 5x + 20 }\) Does that help you?

Parth (parthkohli):

Factor the first two and the last two terms.

OpenStudy (anonymous):

that's it??

OpenStudy (anonymous):

(x+4)*(x+5)

OpenStudy (anonymous):

hmm okay. I knew the 4 and 5 but couldn't figure out how to write them out...

Parth (parthkohli):

Basically you have to express b as the sum of two numbers \(\alpha\) and \(\beta \) such that \(\alpha \times \beta = ac\)

OpenStudy (anonymous):

@ParthKohli I'm sure you know this, but factoring by grouping isn't necessary here? (x + 4)(x + 5) = 0

Parth (parthkohli):

\( \color{Black}{\Rightarrow x( x + 4) + 5(x + 4) }\) \( \color{Black}{\Rightarrow (x + 5)(x + 4) }\) Can you do the rest?

Parth (parthkohli):

@Calcmathlete we have to factor this, rather than solving for x :)

OpenStudy (anonymous):

wow now everyone is confusing me... its all written different

Parth (parthkohli):

May I know where it is written?

Parth (parthkohli):

It's just easy stuff haha you just gotta believe in yourself =)

OpenStudy (anonymous):

its a problem in my text book I am trying to understand how to factor

Parth (parthkohli):

Hmmm how much did you understand?

OpenStudy (anonymous):

the first answer made sense and then the rest was so confusing haha

Parth (parthkohli):

Hmm where did you get confused? I'd be happy to help you :)

OpenStudy (anonymous):

if you could just start from the beginning and re-answer for me that would be fabulous!

OpenStudy (anonymous):

am I done after this: x2+4x+5x+20 ?

Parth (parthkohli):

Okay, so first I'll split the middle terms in two parts: \( \color{Black}{\Rightarrow x^2 + 4x + 5x + 20 }\) Notice how 4 * 5 = a * c

Parth (parthkohli):

Now, you'll have to factor the first two and last two terms.

OpenStudy (anonymous):

I'm not sure how to do that....

Parth (parthkohli):

What is x^2 + 4x? \( \color{Black}{\Rightarrow x^2 + 4x = x(x + 4)}\)

Parth (parthkohli):

Now what is 5x + 20? \( \color{Black}{\Rightarrow 5(x + 4) }\)

Parth (parthkohli):

Now if you write that: \( \color{Black}{\Rightarrow x(x + 4) + 5(x + 4) }\) Now just group them(remember the distributive property?) \( \color{Black}{\Rightarrow (x + 5)(x + 4) }\)

OpenStudy (anonymous):

wait... what happened to the x and the 5 outside ( ) ?

Parth (parthkohli):

Do you remember the distributive property? we're doing the reverse of that.

OpenStudy (anonymous):

can you explain? I'm super rusty on all of this

Parth (parthkohli):

\( \color{Black}{\Rightarrow a(x + n) = ax + an }\) we're essentially reversing that.

OpenStudy (anonymous):

I think I get it

Parth (parthkohli):

Hmm ok

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

@ParthKohli lol. I was just saying before that you can essentially get the same result by finding factors of 20 that add to 9: 4 and 5. Then by pairing it up with an individual x value since the leading coefficient is found by multiplying the first terms in binomials. (x + 4)(x + 5) Factoring by grouping is only necessary if the leading term has a coefficient other than 1. :)

OpenStudy (anonymous):

so the answer is (x + 4)(x + 5) ?

OpenStudy (anonymous):

Yes.